• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name cusanno

    Log entry time 08:03:53 on April18,2010

    Entry number 317978

    keyword=Owl Shift Summary

    00:00 Just restarted production, run 2903
    00:35 start run 2904
    00:59 Dithering monitor fixed
    01:24 start run 2905
    02:15 start run 2906
    03:06 start run 2907
    03:30 Half-Wave-Plate IN
    03:35 start run 2908
    04:25 start run 2909

    04:45 We figure out that the data replay does not work anymore, starting from run 2907.
    After investigation, we understand that the data are OK but the problem is in the replay itself. After several attempts we were able to recover the data replay, simply trying again after a while.

    05:18 start run 2910, it is taken also to understand the problem above
    05:31 start run 2911, now data replay looks working fine again
    06:28 start run 2912

    RC realized that the problems with the replay was related to disk space. He moved the file on another disk and now we are safe for a while but a real fix is needed later on

    07:13 start run 2913
    Shift end with run 2913 still running